OpenStudy (anonymous):

Find g(f(x)) for f(x) = -3x – 6 and G(x)=(-1/2)x+3

OpenStudy (anonymous):

Simplify this (-1/2)(-3x-6)+3

OpenStudy (anonymous):

Multiply (-3x-6) and (-1/2x+3).

OpenStudy (kira_yamato):

Is \[G(x) = \int\limits g(x) dx\]

OpenStudy (anonymous):

radh thats not right and there is no need for integrals here...

OpenStudy (anonymous):

its just simple substitution

OpenStudy (anonymous):

How is it not right? you multiply g(x) * f(x)..?

OpenStudy (anonymous):

3/2x+6 ?

OpenStudy (anonymous):

you dont multiply you substitue f(x) into g(x)

OpenStudy (anonymous):

im saying is this the answer?

OpenStudy (anonymous):

Ah, I understand. Sorry.

OpenStudy (anonymous):

no the answer is 3/2x-15

OpenStudy (anonymous):

oh ok.. thanks!

OpenStudy (anonymous):

What about : Find f(g(x)) for f(x) = -3x – 6 and g(x)=-1/2x+3

OpenStudy (anonymous):

3/2x+6?

OpenStudy (anonymous):

Oh wow im so sorry i switched then up

OpenStudy (anonymous):

3/2x+6 is the right answer to the first one and 3/2x-15 is the answer to the second one

OpenStudy (anonymous):

i see!

OpenStudy (kira_yamato):

I suppose the thing is for all Real values of x. gf(x) = (-1/2)(-3x – 6)+3 gf(x) = 3/2 x + 6

OpenStudy (anonymous):

thanksss

OpenStudy (kira_yamato):

Sorry if I confuse you because I thought G'(x) = g(x) --> saw that in a textbook when I visit Singapore